What was the force being applied by the pitcher's hand if the ball is in contact with it for 0.2s?


1
Expert's answer
2021-03-18T20:11:55-0400

Let's assume that the mass of the ball is "m = 150g = 0.15kg", and the final speed is "v_f = 40m\/s".

Then the accleration of the ball during the contact is:


"a = \dfrac{v_f - v_i}{t}"

where "v_i = 0m\/s" is the initial speed of the ball, and "t = 0.2s" is the time of the contact. Thus, obtain:


"a = \dfrac{40m/s}{0.2s} = 200 m/s^2"

According to the second Newton's law, the force applied by the pitcher's hand is:


"F = ma\\
F = 0.15kg\cdot 200 m/s^2 = 30N"

Answer. 30N.